# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ 4.5 Full Proof of Theorem 1
First of all, the polynomial time complexity follows from the fact that the greedy multi-round policy applies $n\cdot T$ times algorithm Oracle2 to compute the budget vector (via algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$) and $B<n\cdot T$ times the algorithm Oracle1 to select $B$ items, and since both Oracle1 and Oracle2 run in polynomial time. Thus, in the remainder of the proof we will only focus on the approximation guarantee of the greedy multi-round policy.
@@
Fix $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$, where $c\in(0,1/4]$ is a constant that can be arbitrarily fixed. Let $\pi^{*}$ denote an optimal adaptive policy for instance $I$, and consider the following artificial adaptive policy $\hat{\pi}^{*}$ that works in two phases. Let ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$ be a variant of ${\sf BudgetGr}_{\delta,\xi}(I)$ working assuming that $\tilde{\Delta}_{t,i}=\Delta_{t,i}$ holds with probability $1$ (that is, such variant computes the budget vector based on the knowledge of the exact expected increments of the state functions), and let $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ be the resulting budget vector. In the first phase, $\hat{\pi}^{*}$ executes a variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ of the greedy multi-round policy that uses the budget vector $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ to distribute the budget among all $T$ rounds, and let $\vec{S}=(S_{1},\ldots,S_{T})$ denote the aggregated set (with $\sum_{t\in[T]}|S_{t}|=B$) returned by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$. In the second phase, $\pi^{*}$ restarts from round $1$, and by observing the aggregated partial state $\langle\vec{S}\rangle$ it selects further $\min\{nT-B,B\}$ items in an optimal adaptive way over $T$ rounds.
By construction of $\hat{\pi}^{*}$, we can easily observe that the expected performance of policy $\hat{\pi}^{*}$ is not worse than that of $\pi^{*}$, that is
$\sigma(\hat{\pi}^{*})\geq\sigma(\pi^{*}).$ (8)
@@
In the following lemma, by exploiting the adaptive submodularity, we show that the expected value of the aggregated state function achieved in the first phase (that is, $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$) is at least equal to the expected increment of the aggregated state function achieved in the second phase (that is, $\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$), up to an addend $B(\delta+\xi\Lambda)$.
@@
We have $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-2B(\delta+\Lambda\xi)$.
Proof.
Given a round $t\in[T]$ and $j\in[b_{t}^{\prime}]$, let $\bm{v}_{t,i}$ denote the (random) $i$-th item selected at round $t$ by the greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ (when it applies, as sub-routine, the greedy single-round policy) and let $\bm{S}_{t,i}$ denote the (random) set $\{\bm{v}_{t,1},\ldots,\bm{v}_{t,i}\}$. We recall that $\Delta_{t,i}=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ denotes the expected increment of the state function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item at round $t$; furthermore, let $\bm{v}_{t,i}^{*}$ be the (random) item that maximizes $\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)$ and let $\Delta^{*}_{t,i}:=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$.
In the remainder of the proof, we will first show several inequalities (Parts i-iv) and then we will combine them to show the claim (Final Part).
@@
Given a round $t\in[T]$ and $i\in[b_{t}^{\prime}]$, let $A_{t,i}$ denote the event “$\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\leq\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta$”. By the property (i) of Proposition 2 (see Section 4.1.1), we have that event $A_{t,i}$ is true with probability at least $1-\xi$, and then, its complementary event $\overline{A_{t,i}}$ is false with probability at most $\xi$. Thus, we have
@@
Furthermore, given a round $t\in[T]$ and $i,i^{\prime}\in[b_{t}^{\prime}]$ with $i^{\prime}\leq i$, by exploiting the adaptive submodularity we have that:
$\Delta_{t,i^{\prime}}^{*}$ $=\mathbb{E}_{\langle\bm{S}_{t,i^{\prime}-1}\rangle}[\Delta(\bm{v}_{t,i^{\prime}}^{*}|\langle\bm{S}_{t,i^{\prime}-1}\rangle)]$ (10)
$\geq\mathbb{E}_{\langle\bm{S}_{t,i^{\prime}-1}\rangle,\bm{v}_{t,i}^{*}}[\Delta(\bm{v}_{t,i}^{*}|\langle\bm{S}_{t,i^{\prime}-1}\rangle)]$ (11)
$\geq\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle,\bm{v}^{*}_{t,i}}^{*}[\Delta(\bm{v}_{t,i}^{*}|\langle\bm{S}_{t,i-1}\rangle)]$ (12)
$=\Delta^{*}_{t,i},$ (13)
where (11) holds by the optimality of $\bm{v}_{t,i}^{*}$ and (12) holds by adaptive submodularity.
@@
Now, for any $j\in[B]$, let $(t_{j},i_{j})$ be the pair such that $t_{j}$ is the round to which the $j$-th unit of budget is assigned by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$, and $i_{j}$ denotes the budget used at round $t_{j}$ after assigning the $j$-th unit of budget. Given $j\in[B]$, we have that
$\Delta_{t_{B},i_{B}}$ $=\Delta_{t_{B},b^{\prime}_{t_{B}}}$ (14)
$\leq\Delta_{t_{j},b^{\prime}_{t_{j}}}$ (15)
$\leq\Delta^{*}_{t_{j},b^{\prime}_{t_{j}}}$
$\leq\Delta^{*}_{t_{j},i_{j}}$ (16)
$\leq\Delta_{t_{j},i_{j}}+\delta+\xi\Lambda,$ (17)
where (14) holds since $i_{B}$ is equal to the budget assigned at round $t$ at the end of algorithm ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$, (15) holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$, (16) follows from (13) and (17) follows from (9).
By exploiting (17) we get
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$ $=\sum_{j\in[B]}\Delta_{t_{j},i_{j}}$ (18)
$\geq\sum_{j\in[B]}\left(\Delta^{*}_{t_{j},i_{j}}-(\delta+\xi\Lambda)\right)$
$\geq\sum_{j\in[B]}\left(\Delta^{*}_{t_{B},i_{B}}-(\delta+\xi\Lambda)\right)$
$=B\cdot\Delta^{*}_{t_{B},i_{B}}-B(\delta+\xi\Lambda),$
where the first inequality follows from (9) and the second one from (13).
@@
Finally, for any $h\in[\min\{B,Tn-B\}]$, let $\hat{\Delta}^{*}_{h}$ denote the expected increment of the $h$-th item selected by $\hat{\pi}^{*}$ since the beginning of the second phase (in this case, the ordering is not related to the specific round). By construction, we have that $\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}$ is equal to the expected increment achieved in the second phase of $\hat{\pi}^{*}$, that is,
$\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}=\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta}(I)).$ (19)
For a given $h\in[\min\{B,Tn-B\}]$, let $t$ be the round in which the $h$-th item is assigned. We know that $b_{t}$ items have been previously selected in the first phase, when the greedy multi-round policy was run. Thus, by exploiting again the adaptive submodularity, we have that the expected increment $\Delta^{*}_{t,b_{t}+1}$ (i.e., that obtained by selecting optimally an item in round $t$, at the end of the first phase) is at least equal as that achieved by $\hat{\pi}^{*}$ by adding the $h$-th item in the second phase, that is:
$\hat{\Delta}^{*}_{h}\leq\Delta^{*}_{t,b_{t}+1}\leq\Delta_{t,b_{t}+1}+\delta+\xi\Lambda\leq\Delta_{t_{B},i_{B}}+\delta+\xi\Lambda,$ (20)
where the first inequality holds by the above observations, the second one holds by (13) and the last one holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$.
@@
By putting together (18), (19) and (20), we obtain
@@
$\geq B\cdot\Delta_{t_{B},i_{B}}-B(\delta+\xi\Lambda)$
$\geq\min\{B,Tn-B\}\cdot\Delta_{t_{B},i_{B}}-B(\delta+\xi\Lambda)$
@@
$\geq\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}-2B(\delta+\xi\Lambda)$
[…]
and this shows the claim of the lemma. ∎ In the following lemma, we relate the value of ${\sf MultiGr}_{\delta,\xi}(I)$ to that of its variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$.
@@
We recall that $\Delta_{t,i}$ denotes the expected increment of the state function considered in algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, when selecting the $i$-the item at round $t$; by construction of policy ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$, we have that
@@
$\sigma({\sf MultiGr}_{\delta,\xi}(I))=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right].$ (22)
@@
By exploiting similar arguments as in inequality (17) of Lemma 1, one can show that the $\Delta_{t,i}$ is non-increasing in $i$ up to addend $(\delta+\xi\Lambda)$, that is,
@@
for any $t\in[T]$ and $i,i^{\prime}\in[n]$ such that $i^{\prime}\leq i$; in particular, such inequality can be shown by observing that $\Delta_{t,i^{\prime}}+\delta+\xi\Lambda\geq\Delta_{t,i^{\prime}}^{*}\geq\Delta_{t,i}^{*}\geq\Delta_{t,i}$, where $\Delta_{t,j}^{*}$ is the maximum expected increment when adding an item at round $t$, assuming that $j-1$ items have been already selected in that round by the approximate greedy single-round policy (see the proof of Lemma 1 for a formal definition of $\Delta_{t,j}^{*}$).
@@
Now, let $\tilde{\bm{\Delta}}_{t,i}$ be the random quantity considered in algorithm Oracle2 to estimate $\Delta_{t,i}$ with high probability and arbitrarily high precision. Let $A$ be the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$”; by the property (i) of Proposition 3 (see Section 4.1.1), we have that $\mathbb{P}[A]\geq 1-\xi$ and $\mathbb{P}[\overline{A}]=1-\mathbb{P}[A]\leq\xi$, where $\overline{A}$ denotes the complementary event of $A$. Then, given a realization $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true we have that
@@
Let $\overline{\bm{\Delta}}_{t,i}=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$ be the (random) values considered in ${\sf BudgetGr}_{\delta,\xi}(I)$ to assign the budget in a greedy way. Given a realization $(\overline{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\overline{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ of the above values, and the corresponding budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ returned as output by ${\sf BudgetGr}_{\delta,\xi}(I)$, we have that
@@
Indeed, as $\overline{\Delta}_{t,i}$ is non increasing in $i$ (by definition of minimum), one can easily show that the vector $\vec{b}^{*}=(b_{1}^{*},\ldots,b_{T}^{*})$ maximizing $\sum_{i\in[b_{t}^{*}]}\overline{\Delta}_{t,i}$ subject to $\sum_{t\in[T]}b_{t}^{*}=B$ can be computed in a greedy way as in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$, that is, the budget vector $\vec{b}$ returned by ${\sf BudgetGr}_{\delta,\xi}(I)$ is optimal and then (25) holds.
@@
Let $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}$ be a realization of $(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true, and let $\overline{\Delta}_{t,j}:=\min_{i^{\prime}\in[i]}\tilde{\Delta}_{t,i}$. Given $t\in[T],i\in[n]$, let $i^{\prime}\in[i]$ be the lowest integer such that $\tilde{\Delta}_{t,i^{\prime}}=\overline{\Delta}_{t,i}$. By exploiting (23), we get:
@@
$\leq\Delta_{t,i^{\prime}}-\tilde{\Delta}_{t,i^{\prime}}+\delta+\xi\Lambda$ (26)

### v2

@@ 4.5 Full Proof of Theorem 1
Fix $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$, where $c\in(0,1/4]$ is a constant that can be arbitrarily fixed. We first show the polynomial-time complexity of the algorithm. First of all, by Propositions 2 and 3 we have that both Oracle1 and Oracle2 run in polynomial time w.r.t. the size of $I$, $1/\delta$ and $1/\xi$. Then, as $1/\epsilon=O\left(\frac{nT\Lambda}{\lambda\delta}\right)=O\left(\frac{nT\Lambda}{\lambda\xi}\right)$, by the polynomial boundedness property (that polynomially relates quantities $\Lambda$ and $1/\lambda$ to the size of $I$) we conclude that Oracle1 and Oracle2 are polynomial in $1/\epsilon$ and the size of $I$. Finally, the polynomial-time complexity of the whole greedy multi-round policy follows from the fact that such policy applies $n\cdot T$ times algorithm Oracle2 to compute the budget vector (via algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$) and $\sum_{t\in[T]}b_{t}=B<n\cdot T$ times the algorithm Oracle1 to select $b_{t}$ items at each round $t$ (via algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,b_{t})$); the other steps of the greedy multi-round policy are less significant in terms of time-complexity.
@@
Now, we will show the approximation guarantee of the greedy multi-round policy. Let $\pi^{*}$ denote an optimal adaptive policy for instance $I$. Let $\Delta_{t,i}$ be the expected increment of the objective function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item at round $t$, and let $\tilde{\Delta}_{t,i}$ be the estimation of ${\Delta}_{t,i}$ computed in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$ via Oracle2. Let ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$ be a variant of ${\sf BudgetGr}_{\delta,\xi}(I)$ working assuming that the estimation $\tilde{\Delta}_{t,i}$ of $\Delta_{t,i}$ is correct, that is, Oracle2 (differently from Oracle1) is run with $\delta:=\xi:=0$ and then $\tilde{\Delta}_{t,i}=\Delta_{t,i}$ holds with probability $1$. For any fixed $k\in\llbracket 0,B\rrbracket$, let $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ be the budget vector achieved by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$, immediately after the assignment of $k$-th unit of budget. Let $\pi^{\prime}(k)$ denote the variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ of the greedy multi-round policy that uses the budget vector $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ to distribute $k$ units of budget over all $T$ rounds; we observe that ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ coincides with policy $\pi^{\prime}(B)$. Furthermore, consider the following artificial adaptive policy $\hat{\pi}^{*}(k)$ that works in two phases: in the first phase, $\hat{\pi}^{*}(k)$ runs policy $\pi^{\prime}(k)$, so that $k$ items are selected; in the second phase, $\hat{\pi}^{*}(k)$ restarts from round $1$ and, by using also the observations from the first phase, selects further $\min\{nT-k,B\}$ items in an optimal adaptive way over all $T$ rounds.
In the remainder of the proof, we will first relate the value achieved by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ with the optimal adaptive one achieved by $\pi^{*}$, and then we will relate ${\sf MultiGr}_{\delta,\xi}(I)$ with ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ to obtain the desired approximation guarantee.
Fix $k\in\llbracket 0,B-1\rrbracket$. By the monotonicity of the objective functions and since $\hat{\pi}^{*}(k)$ selects as many items as in the optimal adaptive policy $\pi^{*}$, the performance of $\hat{\pi}^{*}(k)$ is not worse than that of $\pi^{*}$, that is,
$\sigma(\hat{\pi}^{*}(k))\geq\sigma(\pi^{*}).$ (12)
@@
In the following lemma, by exploiting the adaptive submodularity, we show that $\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))$ (that is, the expected increment of the objective functions achieved by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$, when selecting the item associated with the $(k+1)$-th unit of budget assigned by ${\sf BudgetGr}^{\prime}_{\delta,\epsilon}(I)$) is at least equal as $\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))$ (that is, the overall expected increment of the aggregated objective function achieved in the second phase of $\sigma(\hat{\pi}^{*}(k))$) divided by $B$, up to an error $\delta+\xi\Lambda$ (due to the sampling procedure of Oracle1).
@@
We have $\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))}{B}-(\delta+\Lambda\xi)$.
Proof of Lemma 1.
Given a round $t\in[T]$ and $i\in[n]$, let $\bm{v}_{t,i}$ denote the (random) $i$-th item selected at round $t$ by the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ and let $\bm{S}_{t,i}$ denote the (random) set $\{\bm{v}_{t,1},\ldots,\bm{v}_{t,i}\}$. We recall that $\Delta_{t,i}=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ is the expected increment of the objective function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item.
In the remainder of the proof, we will show several inequalities (Parts i-iv) that will lead to the claim (Final part).
@@
Given a round $t\in[T]$ and $i\in[n]$, let $\bm{v}_{t,i}^{*}$ be the (random) item that maximizes $\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)$ and $\Delta^{*}_{t,i}:=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$. Furthermore, let $A_{t,i}$ denote the event “$\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\leq\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta$”; by the property (i) of Proposition 2, we have that event $A_{t,i}$ is true with probability at least $1-\xi$, and then, its complementary event $\overline{A_{t,i}}$ is false with probability at most $\xi$. Thus, we have
@@
Now, let $t$ be the round to which the $(k+1)$-th unit of budget is assigned by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$; then, $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ is the budget vector obtained before the above assigned. For any round $q\neq t$, we have that
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))=\Delta_{t,b^{\prime}_{t}(k)+1}$ $\geq\Delta_{q,b^{\prime}_{q}(k)+1}\geq\Delta^{*}_{q,b^{\prime}_{q}(k)+1}-(\delta+\xi\Lambda),$ (14)
where the first inequality holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$ and the second one follows from (13).
@@
For any $h\in[\min\{B,Tn-k\}]$, let $\hat{\Delta}^{*}_{h}(k)$ denote the expected increment of the $h$-th item selected by $\hat{\pi}^{*}(k)$ during the second phase. By construction, we have that $\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)$ is equal to the expected increment achieved in the second phase of $\hat{\pi}^{*}(k)$, that is,
$\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)=\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))).$ (15)
Part iv:
For a given $h\in[\min\{B,Tn-k\}]$, let $\hat{\bm{t}}_{h}$ be the (random) round in which the $h$-th item of the second phase is selected by $\hat{\pi}^{*}(k)$. By exploiting the adaptive submodularity, we can show that $\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}$ (i.e., the expected increment of the objective functions obtained by selecting optimally an item in round $\hat{\bm{t}}_{h}$, at the end of the first phase of $\hat{\pi}^{*}(k)$) is at least equal as $\hat{\Delta}^{*}_{h}(k)$ (i.e., the expected increment of the objective functions achieved by $\hat{\pi}^{*}(k)$ when selecting the $h$-th item during the second phase), independently on the particular realization of $\hat{\bm{t}}_{h}$. Indeed, let $\langle\bm{S}\rangle$ denote the partial state in round $\hat{\bm{t}}_{h}$, observable immediately after the first phase, let $\bm{v}^{*}$ be an item that can be optimally selected in the same round after observing $\langle\bm{S}\rangle$, and let $\langle\hat{\bm{S}}\rangle$ denote the partial state in the same round, but observed immediately before the selection of the $h$-th item $\hat{\bm{v}}^{*}$ in the second phase of $\hat{\pi}^{*}(k)$. We have
$\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}=\mathbb{E}_{\langle\bm{S}\rangle}[\Delta(\bm{v}^{*}|\langle\bm{S}\rangle)]\geq\mathbb{E}_{\langle\hat{\bm{S}}\rangle}[\Delta(\hat{\bm{v}}^{*}|\langle\bm{S}\rangle)]\geq\mathbb{E}_{\langle\hat{\bm{S}}\rangle}[\Delta(\hat{\bm{v}}^{*}|\langle\hat{\bm{S}}\rangle)]=\hat{\Delta}^{*}_{h}(k),$ (16)
where the first inequality holds by the optimality of $\bm{v}^{*}$ and the last one holds since $\langle\bm{S}\rangle\preceq\langle\hat{\bm{S}}\rangle$ and by the adaptive submodularity.
We conclude that
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\Delta^{*}_{\hat{\bm{t}}_{h},b^{\prime}_{\hat{\bm{t}}_{h}}(k)+1}-(\delta+\xi\Lambda)\geq\hat{\Delta}^{*}_{h}(k)-(\delta+\xi\Lambda),$ (17)
where the first inequality holds by (14) (with $q:=\hat{\bm{t}}_{h}$) and the last one holds by (16).
@@
By putting together (15) and (17), we obtain
$B(\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k)))$ $\geq\min\{B,Tn-k\}\cdot(\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k)))$
@@
$\geq\sum_{h\in[\min\{B,Tn-k\}]}\hat{\Delta}^{*}_{h}(k)-B(\delta+\xi\Lambda)$
@@
and, by dividing all members by $B$, the claim of the lemma follows. ∎ By Lemma 1, we obtain
$\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))}{B}-(\delta+\xi\Lambda)\geq\frac{\sigma(\pi^{*})-\sigma(\pi^{\prime}(k))}{B}-(\delta+\xi\Lambda)$
where the last inequality follows from (12). By rearranging the above inequalities we obtain
$\sigma(\pi^{\prime}(k+1))\geq\left(1-\frac{1}{B}\right)\sigma(\pi^{\prime}(k))+\left(\frac{1}{B}\right)\sigma(\pi^{*})-(\delta+\xi\Lambda).$ (18)
By applying recursively (18) from $k=0$ to $k=B-1$ we obtain $\sigma(\pi^{\prime}(k))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}-B(\delta+\xi\Lambda)$. Thus, we have
${\sf MultiGr}^{\prime}_{0,0}(I)$ $=\sigma(\pi^{\prime}(B))$
$\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}-B(\delta+\xi\Lambda)$
$=\left(\frac{1}{B}\right)\sigma(\pi^{*})\left(\frac{1-\left(1-\frac{1}{B}\right)^{B}}{1-\left(1-\frac{1}{B}\right)}\right)-B(\delta+\xi\Lambda)$
$=\left(1-\left(1-\frac{1}{B}\right)^{B}\right)\sigma(\pi^{*})-B(\delta+\xi\Lambda)$
$\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*})-B(\delta+\xi\Lambda),$ (19)
where (19) follows from the known inequality $(1+x)^{r}<e^{rx}$, holding for any $x\geq-1$ and $r>0$, applied with $x=-1/B$ and $r=B$. By (19), we have that ${\sf MultiGr}^{\prime}_{0,0}(I)$ is a $(1-1/e)$-approximate policy.
In the following lemma, we relate the value of ${\sf MultiGr}_{\delta,\xi}(I)$ with that of its variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$; this result, combined with (19), will be used to show the approximation factor of ${\sf MultiGr}_{\delta,\xi}(I)$.
@@
We recall that $\Delta_{t,i}$ denotes the expected increment of the objective function considered in algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, when selecting the $i$-the item at round $t$; by construction of policy ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ and budget vector $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b^{\prime}_{T})$, we have that
@@
$\sigma({\sf MultiGr}_{\delta,\xi}(I))=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right],$ (21)
where the expectation is over the possible realizations of $\vec{\bm{b}}$.
@@
By exploiting similar arguments as in Parts i and iv of Lemma 1, we can show that the $\Delta_{t,i}$ is non-increasing in $i$ up to addend $(\delta+\xi\Lambda)$, that is,
@@
for any $t\in[T]$ and $i,i^{\prime}\in[n]$ such that $i^{\prime}\leq i$. Indeed, let $\Delta_{t,i}^{*}$ be the maximum expected increment obtained by selecting an item at round $t$ in an optimal way, assuming that $i-1$ items have been already selected in that round by policy ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ (see Part i of the proof of Lemma 1 for a formal definition of $\Delta_{t,i}^{*}$). Inequality (22) can be shown by observing that $\Delta_{t,i^{\prime}}+\delta+\xi\Lambda\geq\Delta_{t,i^{\prime}}^{*}\geq\Delta_{t,i}^{*}\geq\Delta_{t,i}$, where the first inequality holds because of the approximate greedy choice of ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ (similarly as in (13)) and the second one holds by the adaptive submodularity (similarly as in (16)).
@@
Now, let $\tilde{\bm{\Delta}}_{t,i}$ be the random quantity considered in algorithm Oracle2 to estimate $\Delta_{t,i}$ with high probability and arbitrarily high precision. Let $A$ be the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$”; by the property (i) of Proposition 3, we have $\mathbb{P}[A]\geq 1-\xi$. Then, given a realization $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true we have that
@@
Let $\overline{\bm{\Delta}}_{t,i}:=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$ be the (random) values considered in ${\sf BudgetGr}_{\delta,\xi}(I)$ to assign the budget in a greedy way. Given a realization $(\overline{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\overline{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ of the above values, and the corresponding budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ returned as output by ${\sf BudgetGr}_{\delta,\xi}(I)$, we have that
@@
Indeed, as $\overline{\Delta}_{t,i}$ is non increasing in $i$ (by definition of minimum), one can easily show that the vector $\vec{b}^{*}=(b_{1}^{*},\ldots,b_{T}^{*})$ maximizing $\sum_{i\in[b_{t}^{*}]}\overline{\Delta}_{t,i}$ subject to $\sum_{t\in[T]}b_{t}^{*}=B$ can be computed in a greedy way as in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$, that is, the budget vector $\vec{b}^{*}$ can be set equal to $\vec{b}$, and then (24) holds by the optimality of $\vec{b}$.
@@
Let $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}$ be a realization of $(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true, and let $\overline{\Delta}_{t,i}$ be the corresponding realization of $\overline{\bm{\Delta}}_{t,i}=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$. Given $t\in[T],i\in[n]$, let $i^{\prime}\in[i]$ be the lowest integer such that $\tilde{\Delta}_{t,i^{\prime}}=\overline{\Delta}_{t,i}$. By exploiting (22), we get:
@@
$\leq\Delta_{t,i^{\prime}}+\delta+\xi\Lambda-\tilde{\Delta}_{t,i^{\prime}}$ (25)